Find an equation of a line that is parallel to the line of 5x-4y=20.
I had to find the x and y intercepts. I have x=5 and y=4. My answer needs to be in y=mx+b form.

Answer by venugopalramana(3286)

Find an equation of a line that is parallel to the line of 5x-4y=20.
I had to find the x and y intercepts. I have x=5 and y=4. My answer needs to be in y=mx+b form.
ANY LINE PARALLEL TO THE ABOVE LINE IS
5X-4Y=K..............WHERE K IS A CONSTANT.
TO FIRM UP THE EQN. WE NEED ANOTHER DATA.
TILL THEN WE CAN ONLY GIVE IN TERMS OF K
X INTERCEPT IS WHEN Y=0,SO 5X=K......X=K/5 IS THE X INTERCEPT.
Y INTERCEPT IS WHEN X=0,SO -4Y=K.....Y=-K/4 IS THE Y INTERCEPT
IF YOU MEANT THAT IT PASSES THROUGH X=5 AND Y=4,THEN WE CAN FIND K
5*5-4*4=K
K=25-16=9
HENCE EQN. IS
5X-4Y=9
THEN X INTERCEPT = K/5=9/5
THEN Y INTERCEPT = -K/4=-9/4